dx/2xz=dy/2yz=dz/z^2-x^2-y^2


1
Expert's answer
2021-09-22T00:01:41-0400

"\frac{dx}{2xz}=\frac{dy}{2yz}=\frac{dz}{z^2-x^2-y^2}"

The first equation gives

"\frac{dx}{x}=\frac{dy}{y}", i.e. "\frac{xdy-ydx}{xy}=\frac{x}{y}d(\frac{y}{x})=0",

"d(\frac{y}{x})=0"

"\frac{y}{x}=k={\rm const}"

The second equation gives

"\frac{dx}{x}=\frac{2zdz}{z^2-x^2-y^2}=\frac{dz^2}{z^2-(1+k^2)x^2}"

"xdz^2-(z^2-(1+k^2)x^2)dx=0"

"\frac{xdz^2-z^2dx}{x^2}-(1+k^2)dx=0"

"d(\frac{z^2}{x}-(1+k^2)x)=0"

"\frac{z^2}{x}-(1+k^2)x=C"

"z^2-(1+k^2)x^2=Cx"

"z^2-x^2-y^2=Cx"

The general integral surface of the given ODE has an equation determined by an arbitrary differentiable function "\Phi" of "(k,C)" :

"\Phi(k,C)=\Phi(\frac{y}{x},\frac{z^2-x^2-y^2}{x})=0"


Answer. "\Phi(\frac{y}{x},\frac{z^2-x^2-y^2}{x})=0", where "\Phi" is an arbitrary differentiable function.